Audrey has 35 dimes and quarters in her piggy bank with a total value of $6.95. How many of the coins are dimes and how many are quarters?
Answer by stanbon(75887)

Audrey has 35 dimes and quarters in her piggy bank with a total value of $6.95. How many of the coins are dimes and how many are quarters?
---
Quantity Eq::: d + q = 35
Value Eq:::::10d + 25q = 695 cents
----------------------------------------
Multiply thru the 1st equation by 10 to get:
10d + 10q = 350
10d + 25q = 695
-------------------
Subtract the 1st equation from the 2nd and solve for "q":
15q = 345
q = 23 (# of quarters)
---
Since d + q = 35, d = 35-23 = 12 (# of dimes)
================================================
Cheers,
Stan H.